Find two functions, f(x) and g(x) such that (fog)(x)=h(x) or (gof)(x)=h(x) where h(x)= sq. rt. of x squared minus 2 over x squared plus 3, and verify.
Answer by stanbon(75887)

Find two functions, f(x) and g(x) such that (fog)(x)=h(x) or (gof)(x)=h(x) where h(x)= sq. rt. of x squared minus 2 over x squared plus 3, and verify.
------------------
fog(x) = sqrt[(x^2-2)/(x^2+3)]
---
Let g(x) = (x^2-2)/(x^2+3)
Let f(x) = sqrt(x)
---
Then fog(x) = f[g(x)] = f[(x^2-2)/(x^2+3)] = sqrt[(x^2-2)/(x^2+3)]
======================
Cheers,
Stan H.